How many grams of CaH2 are needed to generate 147 L of H2 gas if the pressure of H2 is 827 torr at 20 ∘C?
Express your answer using three significant figures.
1
Expert's answer
2018-10-26T08:21:09-0400

CaH2 + 2H2O = Ca(OH)2 + 2H2

n = PV/RT = (827torr)*(147L)/((62.36367L*Torr/K*mol)*(20+273K)) = 6.6531 mol H2

(6.6531mol H2)*(1mol CaH2/2 molH2)*(42.0943 g CaH2/mol) = 140 g CaH2